OpenStudy (anonymous):

The perimeter of a rectangle equals the sum of the lengths of its four sides. If the width of a rectangle is 2 inches and it's perimeter is 30 inches, find the length of the rectangle.

OpenStudy (♪chibiterasu):

Well we have too sides that have a width, so add those two together and subtract the sum from the perimeter: 2^2 = 4 30 - 4 = 26 Now we divide the difference by 2 and: 26/2 = 13 Width = 2 Length = 13

OpenStudy (♪chibiterasu):

two not too

OpenStudy (anonymous):

2 * 2 = 4

OpenStudy (♪chibiterasu):

That is equivalent to 2^2

OpenStudy (♪chibiterasu):

\[2^2\]